# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v4

@@ E.1 Root Cause
For ML-DSA-65:
@@
This matches the NIST ML-DSA specification [FIPS204] (expected 4–5 iterations).
@@ E.2 Throughput Optimizations
  • Parallel Attempts: Generate $B$ independent $(y,w)$ pairs per signing session. Success probability: $1-(0.75)^{B}$.
@@ Appendix F Naive Approach Comparison
Table 13: Improvement over naive threshold approach. “Our Success” values are theoretical estimates based on rejection sampling analysis; measured success rates vary with configuration (see Table 5).
@@ Appendix G Nonce Range Rounding Details
The z-bound check is $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ where $\gamma_{1}-\beta=524092$. Since $524280>524092$, the reduced nonce range does not affect the pass probability. The Irwin-Hall security analysis uses the actual sampled range $N=2\lfloor\gamma_{1}/|S|\rfloor$, so all bounds account for this rounding.
@@ H.2 Hybrid Functionalities
We work in the $(\mathcal{F}_{\mathsf{SPDZ}},\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Rand}},\mathcal{F}_{\mathsf{Broadcast}})$-hybrid model, where:
Functionality H.5 ($\mathcal{F}_{\mathsf{SPDZ}}$ – Authenticated Secret Sharing [DPSZ12]).
@@
Security: $\mathcal{F}_{\mathsf{SPDZ}}$ provides information-theoretic hiding of shares and computational integrity (MAC forgery probability $\leq 1/q$ per check).
Functionality H.6 ($\mathcal{F}_{\mathsf{edaBits}}$ – Extended Doubly-Authenticated Bits [EKMOZ20]).
@@
Functionality H.8 ($\mathcal{F}_{\mathsf{Binary}}$ – Binary Circuit Evaluation).
@@ Appendix I Complete Ideal Functionality for Profile P2
  • $\beta=196$: bound parameter
@@
      Compute pass/fail for each coefficient:
      $\mathsf{pass}_{j}=\begin{cases}1&\text{if }|\tilde{w}^{\prime}_{j}\bmod 2\gamma_{2}|_{[-\gamma_{2},\gamma_{2})}<\mathsf{bound}\\ 0&\text{otherwise}\end{cases}$
      where $|x|_{[-\gamma_{2},\gamma_{2})}$ denotes the centered representative of $x\bmod 2\gamma_{2}$
@@
    Parties compute $\langle\mathsf{masked}_{j}\rangle_{q}=\langle\mathbf{w}^{\prime}_{j}\rangle_{q}+\langle r_{j}\rangle_{q}$ using local addition
@@
    Each party $P_{i}$ computes commitment $\mathsf{Com}_{i,j}=H([\mathsf{masked}_{j}]_{i}\|\mathsf{sid}\|j\|i)$
@@
    Convert public $\mathsf{masked}_{j}$ to binary: $\mathsf{masked}_{j}^{(b)}=(\mathsf{masked}_{j}^{(0)},\ldots,\mathsf{masked}_{j}^{(\lceil\log_{2}q\rceil-1)})$
@@
    For each honest party $P_{i}\in H$:
@@
    Receive corrupted parties’ inputs from $\mathcal{A}$
@@
    Forward corrupted inputs to $\mathcal{F}_{r_{0}\text{-check}}$
@@
    For honest parties: send dummy inputs to $\mathcal{F}_{r_{0}\text{-check}}$ (the functionality only checks if all inputs received)
@@
      Compute simulated $\widetilde{\mathsf{masked}}_{j}=\sum_{i\in H}\tilde{m}_{i,j}+\sum_{i\in C}[\mathsf{masked}_{j}]_{i}\bmod q$
@@
$\mathsf{masked}_{j}=\mathbf{w}^{\prime}_{j}+r_{j}=\sum_{i}[\mathbf{w}^{\prime}]_{i}[j]+\sum_{i}[r_{j}]_{i}\bmod q$
@@
    The public value $\widetilde{\mathsf{masked}}_{j}$ is converted to binary (same as real)
@@
    For binary subtraction circuit: use $\mathcal{F}_{\mathsf{Binary}}$ simulator
@@
    Honest parties’ binary shares are simulated as random (justified by Beaver AND security)
@@
    For each AND gate in the comparison circuit:
    1. (a)
      The gate opens $(d,e)=(x\oplus a,y\oplus b)$ where $(a,b)$ are Beaver triple components
    2. (b)
      Sample $(\tilde{d},\tilde{e})\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$ uniformly
    3. (c)
      This is indistinguishable from real by Lemma L.10
@@
    Simulate all comparison circuits this way
@@
      The AND tree computes $\mathsf{result}=\bigwedge_{j}\mathsf{pass}_{j}$
@@
      Starting from known $\mathsf{result}$, assign intermediate values:
      • If $\mathsf{result}=1$: set all $\mathsf{pass}_{j}=1$
      • If $\mathsf{result}=0$: set at least one $\mathsf{pass}_{j}=0$ (adversary cannot determine which)
    3. (c)
      For each AND gate computing $z=x\land y$:
      • Given target $z$, sample $(x,y)$ consistent with $z$
      • The Beaver triple values $(d,e)$ revealed during computation are uniform regardless of $(x,y)$
@@
The security parameter is $\kappa$ and the MAC key is in $\mathbb{F}_{q}$ where $q\approx 2^{23}$. Per SPDZ security analysis:
  • MAC forgery probability per verification: $1/q\approx 2^{-23}$
  • With $\ell$ verifications: $\epsilon_{\mathsf{SPDZ}}\leq\ell/q$
For our protocol with $m=1536$ coefficients and constant verifications per coefficient:
$\epsilon_{\mathsf{SPDZ}}\leq O(m)/q=O(1536)/8380417<2^{-12}$
@@
For $B=64$ and ML-DSA-65 parameters ($m=nk=1536$ coefficients, $q=8380417$): the cut-and-choose term $2^{-64}$ is negligible, and $\epsilon^{\prime}_{\mathsf{SPDZ}}\leq m/q<2^{-12}$. Thus $\epsilon_{\mathsf{edaBits}}<2^{-64}+2^{-12}<2^{-11.9}$.
@@
where $r_{j}$ is the edaBit value, uniformly distributed in $\mathbb{F}_{q}$ and independent of all other values (by $\mathcal{F}_{\mathsf{edaBits}}$ security).
@@
Proof L.13.
Each AND gate reveals $(d,e)$ which is uniform by Lemma L.10. Replacing real AND computation with $\mathcal{F}_{\mathsf{Binary}}$ (which also produces uniform $(d,e)$) yields identical distributions.
@@
Security implication: These are correctness considerations, not security concerns. The security proof (Lemma above) shows that the adversary’s view—the $(d,e)$ values from Beaver triples—is identically distributed regardless of the underlying $\mathbf{w}^{\prime}_{j}$ values. Correctness ensures the final result $\mathsf{result}$ is computed correctly; security ensures the adversary learns only $\mathsf{result}$.
@@
Case $\mathsf{result}=0$: At least one $\mathsf{pass}_{j}=0$. The simulator must assign leaf values such that the AND tree outputs 0. We use the following procedure:
@@
    Pick a uniformly random index $j^{*}\in[m]$ and set $\mathsf{pass}_{j^{*}}=0$.
@@
    For all $j\neq j^{*}$, set $\mathsf{pass}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}$ uniformly at random.
  3. 3.
    Propagate values through the AND tree (any path containing a 0 produces 0).
This simulation is not identical to the real distribution of leaf values. However, the adversary’s view consists only of Beaver $(d,e)$ pairs, not the leaf values themselves. The following argument shows indistinguishability:
In both cases, the revealed $(d,e)$ values are uniform by Lemma L.10, regardless of the underlying $(x,y)$. The simulator samples uniform $(d,e)\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$ for each AND gate, matching the real distribution exactly.
Key observation: The adversary sees only $(d,e)$ values, never the underlying bits $x,y$ (which remain secret-shared). Since $(d,e)$ is uniform independent of $(x,y)$, backward simulation is perfect.
@@
Proof L.18.
$\mathsf{H}_{5}$ is exactly the execution of simulator $\mathcal{S}$ interacting with $\mathcal{F}_{r_{0}\text{-check}}$:
  • SPDZ replaced with $\mathcal{F}_{\mathsf{SPDZ}}$: done in $\mathsf{H}_{1}$
  • edaBits replaced with $\mathcal{F}_{\mathsf{edaBits}}$: done in $\mathsf{H}_{2}$
  • Masked values simulated as random: done in $\mathsf{H}_{3}$
  • Binary circuits replaced with $\mathcal{F}_{\mathsf{Binary}}$: done in $\mathsf{H}_{4}$
  • AND tree backward-simulated: done in $\mathsf{H}_{5}$
This matches the simulator construction exactly. Thus $\mathsf{H}_{5}\equiv\mathsf{H}_{6}$.
@@
Protocol $\Pi_{r_{0}}$ UC-realizes $\mathcal{F}_{r_{0}\text{-check}}$ in the $(\mathcal{F}_{\mathsf{SPDZ}},\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Rand}},\mathcal{F}_{\mathsf{Broadcast}})$-hybrid model against static malicious adversaries corrupting up to N1N-1 parties, achieving security-with-abort.
@@
$|\Pr[\mathsf{REAL}_{\Pi_{r_{0}},\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}\text{-check}},\mathcal{S},\mathcal{Z}}=1]|\leq\epsilon$
@@
$\epsilon=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$
for ML-DSA-65 parameters.
@@
Concrete bound: For ML-DSA-65 with $q=8380417$ and $m=1536$ coefficients:
  • $\epsilon_{\mathsf{SPDZ}}\leq O(m)/q<2^{-12}$
  • $\epsilon_{\mathsf{edaBits}}<2^{-12}$ (with cut-and-choose parameter $B=64$)
Thus $\epsilon<2^{-11}$, which is negligible in the security parameter.
@@
    Privacy against coalitions of up to N1N-1 parties (by UC security of $\mathcal{F}_{r_{0}\text{-check}}$: only the result bit is leaked)
@@
EUF-CMA: The threshold signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ has identical distribution to single-signer ML-DSA (masks cancel, Lemma 3.4). The security proof of Theorem 4.2 applies directly.

### v5

@@ E.1 Root Cause
For ML-DSA-65 (single-signer approximation: nonce uniform on $\gamma_{1},\gamma_{1}]$):
@@
This matches the NIST ML-DSA specification [FIPS204] (expected 4–5 iterations). Note: In the threshold protocol, nonces follow the Irwin-Hall distribution (sum of $|S|$ uniform variables), so the per-attempt success rates differ from this single-signer estimate; empirical measurements yield 21–45% depending on $T$ and $|S|$ (see Table 13).
@@ E.2 Throughput Optimizations
  • Parallel Attempts: Generate $B$ independent $(y,w)$ pairs per signing session. Success probability: $1-(0.80)^{B}$ (using the theoretical $\approx 20\%$ per-attempt rate; empirical benchmarks yield $\approx 25$$31\%$, giving $1-(0.75)^{B}$ as an optimistic estimate).
@@ Appendix F Naive Approach Comparison
Table 13: Improvement over naive threshold approach. “Our Success” values are empirical measurements from the Rust benchmarks (Table 5). The non-monotone behavior ($32.4\%\to 23.1\%\to 25.0\%$ for $T=16,24,32$) reflects the interplay of two competing effects: (1) smaller per-party nonce range $\gamma_{1}/|S|$ reduces the acceptance probability for the z-bound check; (2) the Irwin-Hall distribution becomes more concentrated near zero for larger $|S|$, partially improving acceptance. The net effect is non-monotone in $|S|$ and configuration-dependent.
@@ Appendix G Nonce Range Rounding Details
The z-bound check is $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ where $\gamma_{1}-\beta=524092$. Since $524280>524092$, the reduced nonce range does not affect the pass probability for this example.
General Proof.
Writing $\gamma_{1}=|S|\cdot\lfloor\gamma_{1}/|S|\rfloor+r$ with $0\leq r\leq|S|-1$, the reduced nonce maximum is $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor=\gamma_{1}-r\geq\gamma_{1}-(|S|-1)$. For the z-bound check to be unaffected, we need $\gamma_{1}-(|S|-1)>\gamma_{1}-\beta$, i.e., $|S|-1<\beta=\tau\eta=196$. This holds for all $|S|\leq 196$, covering all practical signing set sizes. The Irwin-Hall security analysis uses the actual sampled range $N=2\lfloor\gamma_{1}/|S|\rfloor$, so all bounds account for this rounding.
@@ Appendix H UC Security Framework
Remark H.1 (Scope of this appendix).
This appendix provides the extended UC proofs for Profiles P2 and P3+. The Profile P1 UC proof (TEE-based coordinator) is self-contained in Theorem 4.21 and its proof in the main paper; it is not repeated here. This appendix proves UC security for the r0-check subprotocol (Profile P2, Theorem 6.3) and the semi-async signing protocol (Profile P3+, Theorem P.12). These theorems establish UC realization of $\mathcal{F}_{r_{0}\text{-check}}$ and $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ respectively; realization of the full $\mathcal{F}_{\mathsf{ThreshSig}}$ (Definition 4.19) follows by UC composition with the nonce DKG and signing-step sub-protocols.
@@ H.2 Hybrid Functionalities
We work in the $(\mathcal{F}_{\mathsf{SPDZ}},\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Binary}})$-hybrid model, where:
Functionality H.6 ($\mathcal{F}_{\mathsf{SPDZ}}$ – Authenticated Secret Sharing [DPSZ12]).
@@
Security: $\mathcal{F}_{\mathsf{SPDZ}}$ provides statistical hiding of shares (additive shares are perfectly uniform given $\leq N-1$ honest parties) and computational integrity (MAC forgery probability $\leq 1/q$ per check).
Functionality H.7 ($\mathcal{F}_{\mathsf{edaBits}}$ – Extended Doubly-Authenticated Bits [EKMOZ20]).
@@
Functionality H.9 ($\mathcal{F}_{\mathsf{Binary}}$ – Binary Circuit Evaluation).
@@ Appendix I Complete Ideal Functionality for Profile P2
  • $\beta=\tau\cdot\eta=49\cdot 4=196$: bound parameter (ML-DSA-65)
@@
      Compute pass/fail for each coefficient using the FIPS 204 $\mathsf{LowBits}$ function [FIPS204]:
      $\mathsf{pass}_{j}=\begin{cases}1&\text{if }|\mathsf{LowBits}(\mathbf{w}^{\prime}[j],\,2\gamma_{2})|<\mathsf{bound}\\ 0&\text{otherwise}\end{cases}$
      where $\mathsf{LowBits}(r,2\gamma_{2})=r\bmod{\pm}\,2\gamma_{2}$ (the unique representative in $(-\gamma_{2},\gamma_{2}]$ congruent to $r\bmod 2\gamma_{2}$), matching FIPS 204 Algorithm 31. Note: $\mathbf{w}^{\prime}[j]=\sum_{i=1}^{n}\mathsf{received}[i][j]\bmod q$ is first reduced mod $q$, then $\mathsf{LowBits}$ is applied to this $\mathbb{Z}_{q}$ value directly without additional centering mod $q$.
@@
    Parties compute using local addition:
    $\langle\mathsf{masked}_{j}\rangle_{q}=\langle\mathbf{w}^{\prime}_{j}\rangle_{q}+\langle r_{j}\rangle_{q}$
@@
    Each party $P_{i}$ computes commitment:
    $\mathsf{Com}_{i,j}=H\bigl([\mathsf{masked}_{j}]_{i}\|\mathsf{sid}\|j\|i\bigr)$
@@
    Convert public $\mathsf{masked}_{j}$ to binary:
    $\mathsf{masked}_{j}^{(b)}=\bigl(\mathsf{masked}_{j}^{(0)},\ldots,\mathsf{masked}_{j}^{(\lceil\log_{2}q\rceil-1)}\bigr)$
@@
    For each honest party $P_{i}\in H$ (adversary’s view only):
@@
    Receive corrupted parties’ inputs from $\mathcal{A}$; forward corrupted inputs to $\mathcal{F}_{r_{0}\text{-check}}$
@@
    UC framing (honest parties’ inputs): In the ideal execution, honest parties are ideal-world programs that submit their actual inputs directly to $\mathcal{F}_{r_{0}\text{-check}}$ via the ideal-world channel — the simulator $\mathcal{S}$ is not involved in this step. The functionality receives all inputs and computes $\mathsf{result}$; $\mathcal{S}$ receives $\mathsf{result}$ and uses it in Round 8 simulation.
@@
    Consistency: The simulated shares $\tilde{s}_{i,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ are identically distributed to real honest-party shares (SPDZ additive shares are uniformly random given any fixed secret — perfect hiding). No inconsistency arises: the adversary sees random shares, honest parties’ actual values go to $\mathcal{F}$ privately.
@@
      Compute simulated
      $\widetilde{\mathsf{masked}}_{j}=\sum_{i\in H}\tilde{m}_{i,j}+\sum_{i\in C}[\mathsf{masked}_{j}]_{i}\bmod q$
@@
$\mathsf{masked}_{j}$ $=\mathbf{w}^{\prime}_{j}+r_{j}$
$=\sum_{i}[\mathbf{w}^{\prime}]_{i}[j]+\sum_{i}[r_{j}]_{i}\bmod q$
@@
    Convert $\widetilde{\mathsf{masked}}_{j}$ to binary (same as real)
@@
    For binary subtraction: use $\mathcal{F}_{\mathsf{Binary}}$ simulator
@@
    Simulate honest parties’ binary shares as uniform (justified by Beaver AND security)
@@
    For each AND gate, sample $(\tilde{d},\tilde{e})\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$. By Lemma L.11, this matches the real distribution.
@@
    Apply the same simulation to all comparison circuits.
@@
    The adversary’s view of the AND tree is only the Beaver-triple open values $(d_{j},e_{j})$ — the leaf bits $\mathsf{pass}_{j}$ never appear in the adversary’s view.
@@
      For each AND gate $j\in[m]$, sample $(d_{j},e_{j})\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$ uniformly and independently (no leaf assignment is performed).
@@
      Output the simulated view $\{(d_{j},e_{j})\}_{j=1}^{m}$ together with $\mathsf{result}$ received from $\mathcal{F}_{r_{0}\text{-check}}$.
    Justification: By Lemma L.11, the real $(d_{j},e_{j})=(x_{j}\oplus\alpha_{j},y_{j}\oplus\beta_{j})$ is uniform regardless of the inputs $(x_{j},y_{j})$, because $(\alpha_{j},\beta_{j})$ are fresh Beaver-triple components independent of the gate inputs. Since all $m$ Beaver triples are generated independently in preprocessing, the joint distribution is product-uniform. The simulator’s directly-sampled $(d_{j},e_{j})$ therefore matches the real distribution exactly (SD =0=0), for both $\mathsf{result}\in\{0,1\}$ cases.
@@
The security parameter is $\kappa$ and the MAC key is drawn from a dedicated MAC key field $\mathbb{F}_{p}$ with $|p|\geq 64$ bits, independent of the protocol field $\mathbb{F}_{q}$ (this is standard practice in SPDZ implementations [DPSZ12]). Per SPDZ security analysis:
  • MAC forgery probability per verification: $1/p\leq 2^{-64}$
  • With $\ell$ verifications: $\epsilon_{\mathsf{SPDZ}}\leq\ell/p$
For our protocol with $m=1536$ coefficients, MAC verification occurs at OPEN time; the number of verifications is $\ell=m$:
$\epsilon_{\mathsf{SPDZ}}\leq m/p\leq 1536/2^{64}<2^{-54}$
Remark L.4 (Minimal-configuration bound).
If SPDZ is instantiated with the protocol field $\mathbb{F}_{q}$ ($q\approx 2^{23}$) as the MAC key space (minimal configuration), the bound degrades to $\epsilon_{\mathsf{SPDZ}}\leq m/q=1536/8380417<2^{-12.4}$. Practical deployments of P2 must configure SPDZ with a separate MAC field of at least 64 bits to achieve cryptographically negligible distinguishing advantage.
@@
For $B=64$ and ML-DSA-65 parameters ($m=k\cdot N=6\cdot 256=1536$ coefficients, where $N=256$ is the ring degree distinct from party count $n$; $q=8380417$): the cut-and-choose term $2^{-64}$ is negligible, and $\epsilon^{\prime}_{\mathsf{SPDZ}}\leq m/p\leq 1536/2^{64}<2^{-54}$ (using the same 64-bit MAC field; per [EKMOZ20], the SPDZ sub-protocol internal to edaBits opens at most $m$ values per batch). Thus $\epsilon_{\mathsf{edaBits}}<2^{-64}+2^{-54}<2^{-53.9}$.
@@
where $r_{j}$ is the edaBit value, uniformly distributed in $\mathbb{F}_{q}$ (by Remark H.8, rejection sampling ensures $r_{j}$ is exactly uniform) and independent of all other values (by $\mathcal{F}_{\mathsf{edaBits}}$ security).
@@
Proof L.14.
Each AND gate reveals $(d,e)$, which is uniform by Lemma L.11. Replacing real AND computation with $\mathcal{F}_{\mathsf{Binary}}$ (which also produces uniform $(d,e)$) yields identical distributions.
@@
Security implication: These are correctness considerations, not security concerns. The security proof shows that the adversary’s view — the $(d,e)$ values from Beaver triples — is identically distributed regardless of the underlying $\mathbf{w}^{\prime}_{j}$ values. Correctness ensures $\mathsf{result}$ is computed correctly; security ensures the adversary learns only $\mathsf{result}$.
@@
Case $\mathsf{result}=0$: At least one $\mathsf{pass}_{j}=0$. The simulator does not assign leaf values first. Instead, it proceeds directly as follows:
@@
    For each AND gate $j$, sample $(d_{j},e_{j})\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$ uniformly and independently (no reference to any leaf assignment).
@@
    Output the simulated view $\{(d_{j},e_{j})\}_{j=1}^{m}$ along with $\mathsf{result}=0$ from $\mathcal{F}_{r_{0}\text{-check}}$.
No leaf values need be assigned: the leaf bits $(\mathsf{pass}_{j})$ never appear in the adversary’s view — only the $(d_{j},e_{j})$ pairs do. In both cases, the revealed $(d,e)$ values are uniform by Lemma L.11, regardless of the underlying $(x,y)$; the simulator’s sampled $(d_{j},e_{j})$ therefore perfectly matches the real distribution.
This uniformity extends to the joint distribution over the entire AND tree: each gate’s Beaver triple $(\alpha_{j},\beta_{j},\gamma_{j})$ is sampled independently offline; for gate $j$, $(d_{j},e_{j})=(x_{j}\oplus\alpha_{j},y_{j}\oplus\beta_{j})$ is uniform and independent of $(x_{j},y_{j})$ by Lemma L.11. Since triples are pairwise independent across gates, the joint distribution $(d_{1},e_{1},\ldots,d_{m},e_{m})$ is product-uniform, independent of all leaf values.
Key observation: The adversary sees only $(d_{j},e_{j})$ values; the underlying bits $x_{j},y_{j}$ remain secret-shared. Since the joint distribution is product-uniform and independent of the leaf bits, backward simulation is perfect for the entire AND tree.
@@
Proof L.19.
We show that $\mathsf{H}_{5}$ is the ideal execution $\mathsf{H}_{6}$.
After the previous hybrids, every component of the real protocol has been replaced:
  • SPDZ arithmetic shares $\to$ $\mathcal{F}_{\mathsf{SPDZ}}$ outputs (H1)
  • edaBits $\to$ $\mathcal{F}_{\mathsf{edaBits}}$ outputs (H2)
  • Masked reveal values $\to$ uniform random (H3), so adversary’s view of the masked values is simulated
  • Binary comparison circuits $\to$ $\mathcal{F}_{\mathsf{Binary}}$ output (H4)
  • AND tree open values $(d_{j},e_{j})$ $\to$ product-uniform random (H5)
After all these replacements, the only “real” output remaining is the final result bit output by the AND tree. This result bit is computed by $\mathcal{F}_{\mathsf{Binary}}$ (already ideal from H4) and passed to the parties. But $\mathcal{F}_{\mathsf{Binary}}$’s output is exactly the output of $\mathcal{F}_{r_{0}\text{-check}}$: it outputs 1 iff $\|\mathbf{r}_{0}\|_{\infty}<\gamma_{2}-\beta$.
Explicit argument. We verify that the simulator $\mathcal{S}$ from Section K is well-defined and that its output distribution matches $\mathsf{H}_{5}$ component-by-component:
  • SPDZ shares (H1): $\mathcal{S}$ generates random shares $\tilde{s}_{i,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ for honest parties — identical in distribution to $\mathcal{F}_{\mathsf{SPDZ}}$’s output.
  • edaBits (H2): $\mathcal{S}$ uses $\mathcal{F}_{\mathsf{edaBits}}$ outputs — exact match.
  • Masked reveal (H3): $\mathcal{S}$ samples $\tilde{m}_{i,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ — SD =0=0 by one-time-pad argument.
  • Binary comparison (H4): $\mathcal{S}$ uses $\mathcal{F}_{\mathsf{Binary}}$ outputs — exact match.
  • AND-tree pairs (H5): $\mathcal{S}$ samples $(d_{j},e_{j})\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$ — SD =0=0 by Lemma L.11.
  • Result bit: In the ideal execution $\mathsf{H}_{6}$, honest parties submit their actual inputs directly to $\mathcal{F}_{r_{0}\text{-check}}$ via the ideal-world channel (not through $\mathcal{S}$). The functionality computes $\mathsf{result}$ and sends it to $\mathcal{S}$. The simulator completes the adversary view using $\mathsf{result}$.
Each component’s simulated distribution matches $\mathsf{H}_{5}$’s distribution exactly. The result bit is the only component that differs between $\mathsf{H}_{5}$ (computed by $\mathcal{F}_{\mathsf{Binary}}$, which replaced the real AND-tree in $\mathsf{H}_{4}$) and $\mathsf{H}_{6}$ (provided by $\mathcal{F}_{r_{0}\text{-check}}$). But $\mathcal{F}_{\mathsf{Binary}}$’s output equals $\mathcal{F}_{r_{0}\text{-check}}$’s output: both compute $\mathbf{1}[\|\mathbf{r}_{0}\|_{\infty}<\gamma_{2}-\beta]$ on the same inputs. Hence the result-bit distributions are identical, and $\Pr[\mathsf{H}_{5}=1]=\Pr[\mathsf{H}_{6}=1]$.
@@
Protocol $\Pi_{r_{0}}$ UC-realizes $\mathcal{F}_{r_{0}\text{-check}}$ in the hybrid model with $(\mathcal{F}_{\mathsf{SPDZ}},\allowbreak\mathcal{F}_{\mathsf{edaBits}},\allowbreak\mathcal{F}_{\mathsf{Binary}})$ against static malicious adversaries corrupting up to N1N-1 parties, achieving security-with-abort.
@@
$\bigl|\Pr[\mathsf{REAL}_{\Pi_{r_{0}},\mathcal{A},\mathcal{Z}}=1]$ $-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}\text{-check}},\mathcal{S},\mathcal{Z}}=1]\bigr|\leq\epsilon$
@@
$\epsilon=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-53}$
for ML-DSA-65 parameters with a 64-bit dedicated MAC field (recommended configuration; the minimal configuration using the protocol field gives $\epsilon<2^{-11}$).
@@
Concrete bound (recommended configuration, 64-bit MAC field): For ML-DSA-65 with $q=8380417$, $m=1536$ coefficients, and dedicated MAC field $\mathbb{F}_{p}$ with $|p|=64$ bits:
  • $\epsilon_{\mathsf{SPDZ}}\leq m/p=1536/2^{64}<2^{-54}$
  • $\epsilon_{\mathsf{edaBits}}\leq m/p+2^{-64}<2^{-53.9}$ (cut-and-choose $B=64$; $m/p$ is the SPDZ sub-error internal to edaBits)
Thus $\epsilon=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-54}+2^{-53.9}<2^{-53}$, achieving cryptographically negligible distinguishing advantage. Minimal configuration (protocol field as MAC field): $\epsilon_{\mathsf{SPDZ}}<2^{-12.4}$, $\epsilon_{\mathsf{edaBits}}<2^{-11.9}$, total $\epsilon<2^{-11}$; overall signature security relies on Module-SIS at a much higher level (Theorem 4.2).
@@
    r0-check privacy against coalitions of up to N1N-1 parties (by UC security of $\mathcal{F}_{r_{0}\text{-check}}$: only the 1-bit pass/fail result is leaked by the r0-check subprotocol). Note: overall key privacy and unforgeability require $|C|\leq T-1$; if $|C|\geq T$ the adversary can reconstruct the secret key from $T$ Shamir shares. The N1N-1 bound here refers to the r0-check subprotocol’s leakage, not the full signing protocol’s security model.
@@
EUF-CMA: The threshold signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ is syntactically identical in format to single-signer ML-DSA and is accepted by any compliant FIPS 204 verifier (masks cancel in the sum, Lemma 3.6); the response $\mathbf{z}$ follows an Irwin-Hall rather than uniform distribution, with EUF-CMA security loss $0.013\cdot q_{s}$ bits over $q_{s}$ signing queries (per session: <0.013<0.013 bits; Corollary 4.3). The security proof of Theorem 4.2 applies directly.